# Custom Data Augmentation for low resource ASR using Bark and Retrieval-Based Voice Conversion

### v1

@@ 3 Introduction
Significant research efforts are currently dedicated to advancing the realm of Text-to-Speech (TTS) technology. Text-to-speech (TTS) is the task of generating natural-sounding speech given text input. Text-to-speech (TTS) models can be used in any speech-enabled application that requires converting text to speech imitating the human voice. Some of the models are Tortoise TTS, Bark[2], Tachotron[18], Tachotron 2[16], FastSpeech2[12]. Bark is not a conventional text-to-speech model but instead a fully generative text-to-audio model. Bark was developed by Suno which is a transformer-powered text-to-audio system capable of producing exceptionally lifelike speech in multiple languages. Additionally, Bark has the capacity to generate various types of audio content, ranging from music and ambient sounds to basic sound effects. Moreover, the model can create nonverbal expressions such as laughter, sighs, and crying. The methodology presented in this paper focuses on generating a custom corpus similar to Common Voice using small audio clips using Bark, Meta’s Encodec, and HuBert which can be used for personalized tasks and even for custom ASR when dealing with low-resource ASR language like Hindi.
@@ 5.2 RVC model training and inference
The RVC-Project/Retrieval-based-Voice-Conversion-WebUI is an open-source project that allows users to convert voice data into different voices using a retrieval-based voice conversion approach. The project can be found on GitHub[13] and is available for free under the MIT software license. It is recommended that audio recordings be a minimum of 10 minutes long for model training. If the dataset contains noise, it can be removed through Vocals/Accompaniment Separation and Reverbation using the ’HP2-all-vocals’ model. The target sample rate was set to 32,000. The base v2 pre-trained model was utilized for training (f0G32K and f032K) with a batch size of 40 for 200 epochs on the NVIDIA A5000 GPU.

### v2

@@ 3 Introduction
Significant research efforts are currently dedicated to advancing the realm of Text-to-Speech (TTS) technology. Text-to-speech (TTS) is the task of generating natural-sounding speech given text input. Text-to-speech (TTS) models can be used in any speech-enabled application that requires converting text to speech imitating the human voice. Some of the models are Tortoise TTS[3], Bark[2], Tachotron[19], Tachotron 2[17], FastSpeech2[13]. Bark is not a conventional text-to-speech model but instead a fully generative text-to-audio model. Bark was developed by Suno which is a transformer-powered text-to-audio system capable of producing exceptionally lifelike speech in multiple languages. Additionally, Bark has the capacity to generate various types of audio content, ranging from music and ambient sounds to basic sound effects. Moreover, the model can create nonverbal expressions such as laughter, sighs, and crying. The methodology presented in this paper focuses on generating a custom corpus similar to Common Voice using small audio clips using Bark, Meta’s Encodec, and HuBert which can be used for personalized tasks and even for custom ASR when dealing with low-resource ASR language like Hindi.
@@ 5.2 RVC model training and inference
The RVC-Project/Retrieval-based-Voice-Conversion-WebUI is an open-source project that allows users to convert voice data into different voices using a retrieval-based voice conversion approach. The project can be found on GitHub[14] and is available for free under the MIT software license. It is recommended that audio recordings be a minimum of 10 minutes long for model training. If the dataset contains noise, it can be removed through Vocals/Accompaniment Separation and Reverbation using the ’HP2-all-vocals’ model. The target sample rate was set to 32,000. The base v2 pre-trained model was used for training (f0G32K and f032K) with a batch size 40 for 200 epochs on the NVIDIA A5000 GPU. After training for 200 epochs at 40 batch size, the KL Divergence Loss came down to 0.6777.
Figure 1: KL Divergence Loss (loss_kl)